OpenStudy (anonymous):

Find the derivative of f(x) = 6/x at x = -2.

OpenStudy (anonymous):

@zepdrix @SithsAndGiggles @nincompoop @ganeshie8

OpenStudy (anonymous):

What's the derivative of \(\dfrac{6}{x}\)?

OpenStudy (anonymous):

I don't know.....

OpenStudy (anonymous):

Well you have at least three ways to go about it. Power rule, quotient rule, or limit definition. Which of these are you familiar with the most?

OpenStudy (anonymous):

is it 3/2?

OpenStudy (anonymous):

Close, but you're missing a negative sign. Remember that \(\dfrac{1}{x}=x^{-1}\), so the power rule gives \(\dfrac{d}{dx}x^{-1}=(-1)x^{-1-1}=-x^{-2}=-\dfrac{1}{x^2}\).

OpenStudy (anonymous):

so -3/2?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so this one would be 63, right? Find the derivative of f(x) = -2x2 + 11x at x = 9.

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

\(f(x)=-2x^2+11x\) ?

OpenStudy (anonymous):

If so, then I think you made a mistake in your derivative: \(f'(x)=-4x+11\). Or you made a computational error: \(-4(9)+11=-36+11=-25\not=63\).

OpenStudy (anonymous):

ohhhhh okay so it's -25 then, gotcha :D